Thirty-two people were chosen at random from employees of a large company. Their commute times (in hours) were recorded in aConstruct a frequency table using a class interval width of 0.2 starting at 0.15. (Type integers or simplified fractions.) Fr0.95 -1.15 1.15-1.35 1.35 -1.55Which histogram is representative of the above data? A. B. C. Frequency Frequency 16- 12- 8- 4- 16 12- 8- 4- Frequency 16- 12What is the probability that a person chosen at random from the sample will have a commuting time of at least an hour? (Let n
Answer #1
Class interval Frequency Relative Frequency
0.15-0.35 2 0.0625
0.35-0.55 5 0.15625
0.55-0.75 8 0.25
0.75-0.95 6 0.1875
0.95-1.15 6 0.1875
1.15-1.35 4 0.125
1.35-1.55 1 0.03125

Histogram A

P(n\geq 1)=\frac{11}{32}=0.34375

P(n\leq 0.5)=\frac{7}{32}=0.21875